OpenStudy (anonymous):

find the partial der of Q(x,y)=1200x+500y+x^2y-y^2

OpenStudy (psymon):

Well, when doing these, you pick one variable to remain a variable and the other variables to suddenly become constants. So if I first took the derivative with respect to x, I would assume all those y's are suddenly constants. That being said, I would then have 1200 + 2xy. The purely y-terms disappeared due to being constants. Now if I took the partial derivative with respect to y, any x would then be assumed to be a constant. That would give me 500 + x^2 - 2y.

OpenStudy (anonymous):

Qx or Qy?

OpenStudy (anonymous):

Qx = 1200 + 2xy Qy = 500 + x^2 - 2y

OpenStudy (anonymous):